The Soma Inn is trying to determine its break-even point. The inn has 75 rooms that are rented at $ 60 a night. Operating costs are as follows Salaries Utilities Depreciation Maintenance Maid service Other costs $11,150 per month 2,900 per month 1,300 per month 850 per month 8 per room 34 per room Determine the inns break-even point in (1) number of rented rooms per month and (2) dollars. 1. Break-even point in rooms 2. Break-even point per month If the inn plans on renting an average of 50 rooms per day (assuming a 30-day month), what is (1) the monthly margin of safety in dollars and (2) the margin of safety ratio? (Round ratio to 0 decimal places, eg 10%) 1. Margin of safety 2. Margin of safety ratio
Answer #1

Total variable costs=Maid service+Other costs

=(8+34)=$42 per room

Total fixed costs=Salaries+Utilities+Depreciation+Maintenance

=(11150+2900+1300+850)=$16200

Contribution margin=Sales price-Variable cost

=(60-42=$18 per room

Hence breakeven=Fixed cost/Contribution margin

=(16200/18)

=900 rooms

=(900*60)=$54000

2.Margin of safety=Total sales-Breakeven sales

=(50*30)-900=600 rooms

=(600*60)=$36000

Margin of safety ratio=Margin of safetyTotal sales

=(600/(50*30))=40%